# On relative tractor bundles

### v1

@@
Abstract.
This article contributes to the relative BGG-machinery for parabolic geometries. Starting from a relative tractor bundle, this machinery constructs a sequence of differential operators that are naturally associated to the geometry in question. In many situations of interest, it is known that this sequence provides a resolution of a sheaf that can locally be realized as a pullback from a local leaf space of a foliation that is naturally available in this situation. An explicit description of the latter sheaf was only available under much more restrictive assumptions.
For any geometry which admits relative tractor bundles, we construct a large family of such bundles for which we obtain a simple, explicit description of the resolved sheaves under weak assumptions on the torsion of the geometry. In particular, we discuss the cases of Legendrean contact structures and of generalized path geometries, which are among the most important examples for which the relative BGG machinery is available. In both cases, we show that essentially all relative tractor bundles are obtained by our construction and our description of the resolved sheaves applies whenever the BGG sequence is a resolution.
Key words and phrases:
relative BGG-machinery, relative BGG resolution, relative tractor bundle, parabolic geometries, Legendrean contact structure, generalized path geometry
2020 Mathematics Subject Classification:
primary: 58J10; secondary: 53C07, 53C15, 58J60, 58J70
First and second author supported by the Austrian Science Fund (FWF): P33559-N. This article is based upon work from COST Action CaLISTA CA21109 supported by COST (European Cooperation in Science and Technology). https://www.cost.eu.
1. Introduction
@@
The connection to geometry comes from a duality connecting homomorphisms between generalized Verma modules for $(\mathfrak{g},\mathfrak{p})$ to invariant differential operators acting on sections of homogeneous vector bundles over the generalized flag variety $G/P$ (for an appropriate Lie group $G$ with Lie algebra $\mathfrak{g}$ and subgroup $P\subset G$ corresponding to $\mathfrak{p}$) coming from irreducible representations of $P$. The space $G/P$ in turn is the homogeneous model for parabolic geometries of type $(G,P)$, a concept that extends to a general (real or complex) semisimple Lie group $G$ and parabolic subgroup $P\subset G$. The parabolic geometries are defined (uniformly) as Cartan geometries of type $(G,P)$ but there are general results showing that they are equivalent descriptions of simpler underlying structures, whose explicit descriptions are very diverse, see [book]. Among these underlying structures there are well known examples like conformal structures and Levi-non-degenerate CR structures of hypersurface type that are intensively studied with a variety of other methods. In particular, understanding differential operators intrinsic to such geometries has received a lot of interest.
@@
It turned out that things can not only be moved to the geometric side, but the geometric version of the construction then applies to arbitrary curved geometries of type $(G,P)$. Building on [Baston], the first general geometric construction was given in [CSS-BGG], with improvements in [Calderbank-Diemer] and later in [Rel-BGG2]. The basic input for the construction is a so-called tractor bundle that is induced by a representation $\mathbb{V}$ of the group $G$ (which is restricted to $P$). On the homogeneous model $G/P$, any such bundle is canonically trivialized and hence endowed with a canonical linear connection. In curved cases, tractor bundles are non-trivial in general, but they still inherit canonical linear connections, whose curvature equivalently encodes the curvature of the Cartan connection. Coupling these tractor connections to the exterior derivative, one obtains a twisted de Rham sequence.
@@
The BGG construction then compresses the operators in this sequence to a sequence of higher order operators acting on sections of bundles associated to completely reducible representations of $P$, which form the BGG sequence. The latter bundles are induced by Lie algebra homology spaces and Kostant’s version of the Bott–Borel–Weyl theorem [Kostant] implies that they correspond to the Verma modules showing up in Lepowsky’s construction. On locally flat geometries, the de Rham sequence is a fine resolution of the sheaf of local parallel sections of the tractor bundle, and it turns out that the same holds for the induced BGG sequence, so in particular both sequences compute the same cohomology. In the curved case, there still is a close relation between the operators in the BGG sequences and the covariant exterior derivative, which has been exploited very successfully in many applications.
@@
In the setting of the homogeneous models, a relative version of the BGG construction was already used in the book [BEastwood] on the Penrose transform. Here one considers two nested parabolic subgroups $Q\subset P\subset G$, which immediately leads to projection $G/Q\to G/P$. It turns out that the fibers $P/Q$ of this projection again are generalized flag varieties (of a smaller group) and one then uses a “BGG resolution along the fibers”. A geometric version of this relative theory was obtained in [Rel-BGG2] building on algebraic background from [Rel-BGG1]. Starting from a geometry of type $(G,Q)$ on a manifold $M$, the additional parabolic subgroup $P$ gives rise to a natural distribution $T_{\rho}M\subset TM$, called the relative tangent bundle. The second ingredient one needs is a relative tractor bundle which is induced by a completely reducible representation of the group $P$ and the theory of such representations is well understood. It turns out that any such bundle inherits a canonical partial connection (which allows for differentiation in directions in $T_{\rho}M$ only), called the relative tractor connection. With some technical subtleties (in particular in the case that $T_{\rho}M$ is not involutive) an analog of the geometric BGG construction discussed above then produces operators between bundles associated to completely reducible representations of $Q$.
@@
For appropriate initial representations, the relative BGG construction produces operators that cannot be obtained by the original BGG construction, in particular in cases of so-called singular infinitesimal character. In other cases, it provides alternative constructions of BGG operators. A very important feature is that the relative construction provides complexes and resolutions in many non-flat cases. On the one hand, this needs involutivity of the distribution $T_{\rho}M$, which can be easily characterized in terms of the curvature and the harmonic curvature of the original geometry of type $(G,Q)$. Under slightly stronger curvature conditions one obtains a fine resolution of a sheaf that locally is a pull-back of a sheaf on a local leaf space for the distribution $T_{\rho}M$. An issue that remained open in [Rel-BGG2] is how to concretely identify this sheaf, even in the case that $M$ is globally the total space of a bundle over some base space $N$ with vertical subbundle $T_{\rho}M$.
@@
It is exactly this point that we address in the current paper. We show that for each case $Q\subset P\subset G$, there is a large class of relative tractor bundles which, assuming involutivity of $T_{\rho}M$ and a weak condition on the torsion of the geometry, can locally be explicitly identified with pullbacks of vector bundles on appropriate local leaf spaces for $T_{\rho}M$. Under a slightly stronger condition on the torsion, the sections of these relative tractor bundles, which are parallel for the relative tractor connection, are shown to locally be exactly the pullbacks of sections of the bundles on local leaf spaces. These are exactly the sheaves that, assuming appropriate conditions on the curvature, are resolved by the corresponding relative BGG sequence. Moreover, we can explicitly relate the relative tractor connections on this bundle to well known operations, like the adjoint tractor connection and (in the case that $T_{\rho}M$ is involutive) to the Bott connection associated to the induced foliation. These results originally arose during the work on the PhD theses of the second and third author under the direction of the first author for two specific examples of structures.
@@
It should be remarked at this point that the relative theory makes sense only if $Q$ does not correspond to a maximal parabolic subalgebra of $\mathfrak{g}$, since otherwise there is no intermediate parabolic subgroup $P$. Consequently, the theory does not apply to examples like conformal structures or CR structures of hypersurface type. Still, there are at least two important examples of parabolic geometries for which relative BGG sequences are available, namely Legendrean (or Lagrangean) contact structures and generalized path geometries and these are the two examples referred to above. Both these structures can exist only on manifolds of odd dimension. The former consists of a contact structure together with a chosen decomposition of the contact subbundle into the direct sum of two Legendrean subbundles. These structures are closely similar to CR structures and have been studied intensively during the last years and successfully applied to classification problems in CR geometry, see e.g. [Doubrov-Medvedev-The], [Doubrov-Merker-The] and [MFMZ]. For this geometry, our construction essentially provides all relative tractor bundles. Generalized path geometries are also defined by a configuration of distributions, see §3.5 below. Their importance comes from the fact that they provide an equivalent encoding of systems of second order ODEs, see [Fels]. For generalized path geometries, we obtain essentially all relative tractor bundles for one of the possible intermediate parabolic subgroups and a large subclass for the other.
@@
Let us briefly outline the contents of the paper. In §2, we first collect the necessary setup for nested parabolics. As a new ingredient compared to the discussion in [Rel-BGG1, Rel-BGG2] we focus on the natural bigrading of the Lie algebra $\mathfrak{g}$ induced by the nested parabolic subalgebras $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$, which is very convenient for our purposes. Considering an appropriate filtration induced by this bigrading, one quickly arrives at a class of representations that induce basic relative tractor bundles, which can be obtained as subquotients of the adjoint tractor bundle of the geometry.
@@
In §3, we start by showing that the relative tractor connections on the bundles in our class are induced by the adjoint tractor connection. Next we derive an explicit formula for the relative tractor connections in terms of the Lie bracket of vector fields and the torsion of the geometry. The relation to the Lie bracket is crucial for the proof of Theorem 8, which is the main technical result of our article. Under appropriate conditions on the torsion, we obtain a filtration of the tangent bundle of local leaf spaces and relate the relative tractor bundles we have constructed to the components of the associated graded vector bundle to this filtration. Under slightly stronger conditions, it is shown that sections, which are parallel for the relative tractor connection, coincide with pullbacks of sections on the local leaf space. This quickly leads to a large class of relative tractor bundles for which similar results are available, see §3.3.
@@
In the last part of the article we discuss the two examples which gave rise to the developments of this article, namely Legendrean contact structures in §3.4 and (generalized) path geometries in §3.5. In both cases the situation is relatively simple in the sense that the filtrations of the tangent bundle of local leaf spaces are trivial, so the special relative tractor bundles we construct are all directly related to tensor bundles on local leaf spaces. For both examples of structures, there are two possible intermediate parabolic subgroups between $Q$ and $G$. We in particular prove that, up to isomorphism, our construction produces essentially all relative tractor bundles for both intermediate groups in the Legendrean contact case and for one of the intermediate groups in the case of generalized path geometries.
2. A basic class of relative tractor bundles
@@
We start by briefly recalling the setup of two nested parabolic subalgebras $\mathfrak{q}\subset\mathfrak{p}$ in a semisimple Lie algebra $\mathfrak{g}$ with a compatible choice $Q\subset P\subset G$ of groups as discussed in [Rel-BGG1]. The focus here is on a natural bigrading induced by the pair $(\mathfrak{q},\mathfrak{p})$.
2.1. The bigrading determined by a pair of parabolics
@@
There is a well known correspondence between standard parabolic subalgebras of a real or complex semisimple Lie algebra $\mathfrak{g}$, gradings of $\mathfrak{g}$, and (after an appropriate choice of Cartan subalgebra) subsets of the set of simple roots respectively simple restricted roots. By definition, any (restricted) root can be uniquely written as a linear combination of simple (restricted) roots with integer coefficients that are either all $\geq 0$ or all $\leq 0$. Given a subset $\Sigma$ of simple (restricted) roots, one defines the $\Sigma$-height of a root $\alpha$ to be the sum of the coefficients of all elements of $\Sigma$ in this expansion of $\alpha$. The grading corresponding to $\Sigma$ is then defined by putting the Cartan subalgebra into degree zero and giving (restricted) root spaces degree equal to the $\Sigma$-height of the root. The compatibility of the (restricted) root decomposition with the Lie bracket of $\mathfrak{g}$, readily implies compatibility of this grading with the Lie bracket. In particular, the sum of all grading components of non-positive degree is a subalgebra of $\mathfrak{g}$ and this is the parabolic subalgebra determined by $\Sigma$. See Sections 3.2.1 and 3.2.9 of [book] for details.
@@
Definition 1.
Consider a real or complex semisimple Lie algebra $\mathfrak{g}$ endowed with two nested standard parabolic subalgebras $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$ corresponding to subsets $\Sigma_{\mathfrak{p}}\subset\Sigma_{\mathfrak{q}}$ of simple (restricted) roots. Then we define $\mathfrak{g}_{(0,0)}\subset\mathfrak{g}$ to be the sum of the Cartan subalgebra and all (restricted) root spaces of roots for which both the $\Sigma_{\mathfrak{p}}$-height and the $\Sigma_{\mathfrak{q}}$-height is zero. For integers $i^{\prime},i^{\prime\prime}$ which have the same sign and are not both zero, we define $\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ to be the direct sum of all the restricted root spaces for roots of $\Sigma_{\mathfrak{p}}$-height $i^{\prime}$ and $\Sigma_{\mathfrak{q}}$-height $i^{\prime}+i^{\prime\prime}$.
By construction we have a decomposition $\mathfrak{g}=\oplus_{i^{\prime},i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ and we start by clarifying some elementary properties. Recall that the nilradical $\mathfrak{p}_{+}$ of a parabolic $\mathfrak{p}$ is the sum all positive grading components, while the reductive Levi-factor $\mathfrak{p}_{0}$ is the degree zero component of the grading determined by $\mathfrak{p}$.
Proposition 2.
The decomposition $\mathfrak{g}=\oplus_{i^{\prime},i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ induced by $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$ has the following properties.
(i) $[\mathfrak{g}_{(i^{\prime},i^{\prime\prime})},\mathfrak{g}_{(j^{\prime},j^{\prime\prime})}]\subset\mathfrak{g}_{(i^{\prime}+i^{\prime\prime},j^{\prime}+j^{\prime\prime})}$, so one obtains a bigrading of the Lie algebra $\mathfrak{g}$. In particular, the bracket vanishes on $\mathfrak{g}_{(0,i^{\prime\prime})}\times\mathfrak{g}_{(i^{\prime},0)}$ if $i^{\prime}>0$ and $i^{\prime\prime}<0$ or $i^{\prime}<0$ and $i^{\prime\prime}>0$.
(ii) $\mathfrak{p}=\oplus_{i^{\prime}\geq 0,i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$, $\mathfrak{p}_{+}=\oplus_{i^{\prime}>0,i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$, and $\mathfrak{p}_{0}=\oplus_{i^{\prime\prime}}\mathfrak{g}_{(0,i^{\prime\prime})}$.
(iii) $\mathfrak{q}=\oplus_{i^{\prime},i^{\prime\prime}\geq 0}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ and $\mathfrak{q}_{+}=\oplus_{i^{\prime}+i^{\prime\prime}>0}g_{(i^{\prime},i^{\prime\prime})}$, and $\mathfrak{q}_{0}=\mathfrak{g}_{(0,0)}$.
@@
The decomposition we obtain is a refinement of the one in formula (2.1) of [Rel-BGG1]. In the notation used there, we get $\mathfrak{p}_{-}=\oplus_{i^{\prime}<0,i^{\prime\prime}\leq 0}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$, $\mathfrak{p}_{0}\cap\mathfrak{q}_{-}=\oplus_{i^{\prime\prime}<0}\mathfrak{g}_{(0,i^{\prime\prime})}$, $\mathfrak{p}_{0}\cap\mathfrak{q}_{+}=\oplus_{i^{\prime\prime}>0}\mathfrak{g}_{(0,i^{\prime\prime})}$ for the components are not determined in Proposition 2.1.
2.2. The group level and invariant filtrations

### v2

@@
[…]
On Relative Tractor Bundles
[…]
On Relative Tractor Bundles
[…]
Andreas ČAP, Zhangwen GUO and Michał Andrzej WASILEWICZ
[…]
A. Čap, Z. Guo and M.A. Wasilewicz
[…]
Received June 03, 2024, in final form November 18, 2024; Published online November 29, 2024
[…]
This article contributes to the relative BGG-machinery for parabolic geometries. Starting from a relative tractor bundle, this machinery constructs a sequence of differential operators that are naturally associated to the geometry in question. In many situations of interest, it is known that this sequence provides a resolution of a sheaf that can locally be realized as a pullback from a local leaf space of a foliation that is naturally available in this situation. An explicit description of the latter sheaf was only available under much more restrictive assumptions. For any geometry which admits relative tractor bundles, we construct a large family of such bundles for which we obtain a simple, explicit description of the resolved sheaves under weak assumptions on the torsion of the geometry. In particular, we discuss the cases of Legendrean contact structures and of generalized path geometries, which are among the most important examples for which the relative BGG machinery is available. In both cases, we show that essentially all relative tractor bundles are obtained by our construction and our description of the resolved sheaves applies whenever the BGG sequence is a resolution.
[…]
relative BGG-machinery; relative BGG resolution; relative tractor bundle; parabolic geometries; Legendrean contact structure; generalized path geometry
[…]
58J10; 53C07; 53C15; 58J60; 58J70
1 Introduction
Tractor bundles and relative tractor bundles are a special class of vector bundles that are naturally associated to manifolds endowed with a geometric structure from the class of parabolic geometries. A major reason for their importance lies in the role in the construction of Bernstein–Gelfand–Gelfand sequences (or BGG sequences) and in the relative version of this construction, respectively. To explain this, we have to review some background: The motivation for these constructions comes from algebraic results in the realm of semisimple Lie algebras. The starting point was [3], in which of I.N. Bernšteĭn, I.M. Gel’fand and S.I. Gel’fand constructed a resolution of any finite-dimensional representation of a complex semisimple Lie algebra $\mathfrak{g}$ by homomorphisms of Verma modules. This was generalized by J. Lepowsky to resolutions by homomorphisms of generalized Verma modules for a given parabolic subalgebra $\mathfrak{p}\subset\mathfrak{g}$, see [17].
@@ 1 Introduction
The connection to geometry comes from a duality connecting homomorphisms between generalized Verma modules for $(\mathfrak{g},\mathfrak{p})$ to invariant differential operators acting on sections of homogeneous vector bundles over the generalized flag variety $G/P$ (for an appropriate Lie group $G$ with Lie algebra $\mathfrak{g}$ and subgroup $P\subset G$ corresponding to $\mathfrak{p}$) coming from irreducible representations of $P$. The space $G/P$ in turn is the homogeneous model for parabolic geometries of type $(G,P)$, a concept that extends to a general (real or complex) semisimple Lie group $G$ and parabolic subgroup $P\subset G$. The parabolic geometries are defined (uniformly) as Cartan geometries of type $(G,P)$ but there are general results showing that they are equivalent descriptions of simpler underlying structures, whose explicit descriptions are very diverse, see [7]. Among these underlying structures there are well-known examples like conformal structures and Levi-non-degenerate CR structures of hypersurface type that are intensively studied using a variety of other methods. In particular, understanding differential operators intrinsic to such geometries has received a lot of interest.
@@
It turned out that things can not only be moved to the geometric side, but the geometric version of the construction then applies to arbitrary curved geometries of type $(G,P)$. Building on [1], the first general geometric construction was given in [8], with improvements in [4] and later in [10]. The basic input for the construction is a so-called tractor bundle that is induced by a representation $\mathbb{V}$ of the group $G$ (which is restricted to $P$). On the homogeneous model $G/P$, any such bundle is canonically trivialized and hence endowed with a canonical linear connection. In curved cases, tractor bundles are non-trivial in general, but they still inherit canonical linear connections, whose curvature equivalently encodes the curvature of the Cartan connection. Coupling these tractor connections to the exterior derivative, one obtains a twisted de Rham sequence.
@@
The BGG construction then compresses the operators in this sequence to a sequence of higher order operators acting on sections of bundles associated to completely reducible representations of $P$, which form the BGG sequence. The latter bundles are induced by Lie algebra homology spaces and Kostant’s version of the Bott–Borel–Weyl theorem [16] implies that they correspond to the Verma modules showing up in Lepowsky’s construction. On locally flat geometries, the de Rham sequence is a fine resolution of the sheaf of local parallel sections of the tractor bundle, and it turns out that the same holds for the induced BGG sequence, so in particular both sequences compute the same cohomology. In the curved case, there still is a close relation between the operators in the BGG sequences and the covariant exterior derivative, which has been exploited very successfully in many applications.
@@
In the setting of the homogeneous models, a relative version of the BGG construction was already used in the book [2] on the Penrose transform. Here one considers two nested parabolic subgroups $Q\subset P\subset G$, which immediately leads to a projection $G/Q\to G/P$. It turns out that the fibers $P/Q$ of this projection again are generalized flag varieties (of a smaller group) and one then uses a “BGG resolution along the fibers”. A geometric version of this relative theory was obtained in [10] building on algebraic background from [9]. Starting from a geometry of type $(G,Q)$ on a manifold $M$, the additional parabolic subgroup $P$ gives rise to a natural distribution ${T_{\rho}M\subset TM}$, called the relative tangent bundle. The second ingredient one needs is a relative tractor bundle which is induced by a completely reducible representation of the group $P$ and the theory of such representations is well understood. It turns out that any such bundle inherits a canonical partial connection (which allows for differentiation in directions in $T_{\rho}M$ only), called the relative tractor connection. With some technical subtleties (in particular in the case that $T_{\rho}M$ is not involutive) an analog of the geometric BGG construction discussed above then produces operators between bundles associated to completely reducible representations of $Q$.
@@
For appropriate initial representations, the relative BGG construction produces operators that cannot be obtained by the original BGG construction, in particular in cases of so-called singular infinitesimal character. In other cases, it provides alternative constructions of BGG operators. A very important feature is that the relative construction provides complexes and resolutions in many non-flat cases. On the one hand, this needs involutivity of the distribution $T_{\rho}M$, which can be easily characterized in terms of the curvature and the harmonic curvature of the original geometry of type $(G,Q)$. Under slightly stronger curvature conditions one obtains a fine resolution of a sheaf that locally is a pull-back of a sheaf on a local leaf space for the distribution $T_{\rho}M$. An issue that remained open in [10] is how to concretely identify this sheaf, even in the case that $M$ is globally the total space of a bundle over some base space $N$ with vertical subbundle $T_{\rho}M$.
@@
It is exactly this point that we address in the current paper. Given $Q\subset P\subset G$ as above, a geometry of type $(G,Q)$ on $M$ canonically determines natural subbundles $T^{i^{\prime}}_{P}M\subset TM$, see Section 2.3. The conditions on the torsion of the geometry referred to below are on its behavior with respect to these subbundles. We show in each such situation, there is a large class of relative tractor bundles which, assuming involutivity of $T_{\rho}M$ and a weak condition on the torsion, can locally be explicitly identified with pullbacks of vector bundles on appropriate local leaf spaces for $T_{\rho}M$. This is proved in part (1) of Theorem 3.4. Part (2) of this theorem shows that under a slightly stronger condition on the torsion, the sections of these relative tractor bundles, which are parallel for the relative tractor connection, locally are exactly the pullbacks of sections of the bundles on local leaf spaces. These are exactly the sheaves that, assuming appropriate conditions on the curvature, are resolved by the corresponding relative BGG sequence. Moreover, we can explicitly relate the relative tractor connections on this bundle to well-known operations, like the adjoint tractor connection and (in the case that $T_{\rho}M$ is involutive) to the Bott connection associated to the induced foliation, see Theorem 3.3. These results originally arose during the work on the PhD theses of the second and third author under the direction of the first author for two specific examples of structures.
@@
It should be remarked at this point that the relative theory makes sense only if $Q$ does not correspond to a maximal parabolic subalgebra of $\mathfrak{g}$, since otherwise there is no intermediate parabolic subgroup $P$. Consequently, the theory does not apply to examples like conformal structures or CR structures of hypersurface type. Still, there are at least two important examples of parabolic geometries for which relative BGG sequences are available, namely Legendrean (or Lagrangean) contact structures and generalized path geometries and these are the two examples referred to above. Both these structures can exist only on manifolds of odd dimension. The former consists of a contact structure together with a chosen decomposition of the contact subbundle into the direct sum of two Legendrean subbundles. These structures are closely similar to CR structures and have been studied intensively during the last years and successfully applied to classification problems in CR geometry, see, e.g., [11, 12] and [18]. For this geometry, our construction essentially provides all relative tractor bundles. Generalized path geometries are also defined by a configuration of distributions, see Section 3.5 below. Their importance comes from the fact that they provide an equivalent encoding of systems of second order ODEs, see [14]. For generalized path geometries, we obtain essentially all relative tractor bundles for one of the possible intermediate parabolic subgroups and a large subclass for the other.
@@
Let us briefly outline the contents of the paper. In Section 2, we first collect the necessary setup for nested parabolics. As a new ingredient compared to the discussion in [9, 10], we focus on the natural bigrading of the Lie algebra $\mathfrak{g}$ induced by the nested parabolic subalgebras $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$, which is very convenient for our purposes. Considering an appropriate filtration induced by this bigrading, one quickly arrives at a class of representations that induce basic relative tractor bundles, which can be obtained as subquotients of the adjoint tractor bundle of the geometry.
@@
In Section 3, we start by showing that the relative tractor connections on the bundles in our class are induced by the adjoint tractor connection. Next, we derive an explicit formula for the relative tractor connections in terms of the Lie bracket of vector fields and the torsion of the geometry. The relation to the Lie bracket is crucial for the proof of Theorem 3.4, which is the main technical result of our article. Under appropriate conditions on the torsion, we obtain a filtration of the tangent bundle of local leaf spaces and relate the relative tractor bundles we have constructed to the components of the associated graded vector bundle to this filtration. Under slightly stronger conditions, it is shown that sections, which are parallel for the relative tractor connection, coincide with pullbacks of sections on the local leaf space. This quickly leads to a large class of relative tractor bundles for which similar results are available, see Section 3.3.
@@
In the last part of the article, we discuss the two examples which gave rise to the developments of this article, namely Legendrean contact structures in Section 3.4 and (generalized) path geometries in Section 3.5. In both cases, the situation is relatively simple in the sense that the filtrations of the tangent bundle of local leaf spaces are trivial, so the special relative tractor bundles we construct are all directly related to tensor bundles on local leaf spaces. For both examples of structures, there are two possible intermediate parabolic subgroups between $Q$ and $G$. We in particular prove that, up to isomorphism, our construction produces essentially all relative tractor bundles for both intermediate groups in the Legendrean contact case and for one of the intermediate groups in the case of generalized path geometries.
2 A basic class of relative tractor bundles
@@
We start by briefly recalling the setup of two nested parabolic subalgebras $\mathfrak{q}\subset\mathfrak{p}$ in a semisimple Lie algebra $\mathfrak{g}$ with a compatible choice $Q\subset P\subset G$ of groups as discussed in [9]. The focus here is on a natural bigrading induced by the pair $(\mathfrak{q},\mathfrak{p})$.
2.1 The bigrading determined by a pair of parabolics
@@
There is a well-known correspondence between standard parabolic subalgebras of a real or complex semisimple Lie algebra $\mathfrak{g}$, gradings of $\mathfrak{g}$, and (after an appropriate choice of Cartan subalgebra) subsets of the set of simple roots respectively simple restricted roots. By definition, any (restricted) root can be uniquely written as a linear combination of simple (restricted) roots with integer coefficients that are either all $\geq 0$ or all $\leq 0$. Given a subset $\Sigma$ of simple (restricted) roots, one defines the $\Sigma$-height of a root $\alpha$ to be the sum of the coefficients of all elements of $\Sigma$ in this expansion of $\alpha$. The grading corresponding to $\Sigma$ is then defined by putting the Cartan subalgebra into degree zero and giving (restricted) root spaces degree equal to the $\Sigma$-height of the root. The compatibility of the (restricted) root decomposition with the Lie bracket of $\mathfrak{g}$, readily implies compatibility of this grading with the Lie bracket. In particular, the sum of all grading components of non-positive degree is a subalgebra of $\mathfrak{g}$ and this is the parabolic subalgebra determined by $\Sigma$. See [7, Sections 3.2.1 and 3.2.9] for details.
@@
Definition 2.1.
Consider a real or complex semisimple Lie algebra $\mathfrak{g}$ endowed with two nested standard parabolic subalgebras $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$ corresponding to subsets $\Sigma_{\mathfrak{p}}\subset\Sigma_{\mathfrak{q}}$ of simple (restricted) roots. Then we define $\mathfrak{g}_{(0,0)}\subset\mathfrak{g}$ to be the sum of the Cartan subalgebra and all (restricted) root spaces of roots for which both the $\Sigma_{\mathfrak{p}}$-height and the $\Sigma_{\mathfrak{q}}$-height is zero. For integers $i^{\prime}$$i^{\prime\prime}$ which have the same sign and are not both zero, we define $\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ to be the direct sum of all the restricted root spaces for roots of $\Sigma_{\mathfrak{p}}$-height $i^{\prime}$ and $\Sigma_{\mathfrak{q}}$-height $i^{\prime}+i^{\prime\prime}$.
We remark that the set $\Sigma_{\mathfrak{q}}\setminus\Sigma_{\mathfrak{p}}$ of simple roots gives rise to a second parabolic subalgebra $\tilde{\mathfrak{p}}\subset\mathfrak{g}$ such that $\mathfrak{q}=\mathfrak{p}\cap\tilde{\mathfrak{p}}$. The integers $i^{\prime\prime}$ in Definition 2.1 correspond to the grading of $\mathfrak{g}$ induced by that parabolic. Since $\tilde{\mathfrak{p}}$ does not play a role in the geometric developments below, we avoid using it explicitly.
By construction we have a decomposition $\mathfrak{g}=\bigoplus_{i^{\prime},i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ and we start by clarifying some elementary properties. Recall that the nilradical $\mathfrak{p}_{+}$ of a parabolic $\mathfrak{p}$ is the sum all positive grading components, while the reductive Levi-factor $\mathfrak{p}_{0}$ is the degree zero component of the grading determined by $\mathfrak{p}$.
Proposition 2.2.
The decomposition $\mathfrak{g}=\bigoplus_{i^{\prime},i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ induced by $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$ has the following properties:
  • $(i)$
    $\bigl{[}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})},\mathfrak{g}_{(j^{\prime},j^{\prime\prime})}\bigr{]}\subset\mathfrak{g}_{(i^{\prime}+i^{\prime\prime},j^{\prime}+j^{\prime\prime})}$, so one obtains a bigrading of the Lie algebra $\mathfrak{g}$. In particular, the bracket vanishes on $\mathfrak{g}_{(0,i^{\prime\prime})}\times\mathfrak{g}_{(i^{\prime},0)}$ if $i^{\prime}>0$ and $i^{\prime\prime}<0$ or $i^{\prime}<0$ and $i^{\prime\prime}>0$.
  • $(ii)$
    $\mathfrak{p}=\bigoplus_{i^{\prime}\geq 0,i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$, $\mathfrak{p}_{+}=\bigoplus_{i^{\prime}>0,i^{\prime\prime}}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$, and $\mathfrak{p}_{0}=\bigoplus_{i^{\prime\prime}}\mathfrak{g}_{(0,i^{\prime\prime})}$.
  • $(iii)$
    $\mathfrak{q}=\bigoplus_{i^{\prime},i^{\prime\prime}\geq 0}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$ and $\mathfrak{q}_{+}=\bigoplus_{i^{\prime}+i^{\prime\prime}>0}g_{(i^{\prime},i^{\prime\prime})}$, and $\mathfrak{q}_{0}=\mathfrak{g}_{(0,0)}$.
@@
The decomposition we obtain is a refinement of the one in formula (2.1) of [9]. In the notation used there, we get $\mathfrak{p}_{-}=\bigoplus_{i^{\prime}<0,i^{\prime\prime}\leq 0}\mathfrak{g}_{(i^{\prime},i^{\prime\prime})}$, $\mathfrak{p}_{0}\cap\mathfrak{q}_{-}=\bigoplus_{i^{\prime\prime}<0}\mathfrak{g}_{(0,i^{\prime\prime})}$, $\mathfrak{p}_{0}\cap\mathfrak{q}_{+}=\bigoplus_{i^{\prime\prime}>0}\mathfrak{g}_{(0,i^{\prime\prime})}$ for the components are not determined in Proposition 2.1.
2.2 The group level and invariant filtrations